3. A large board L = 6.00 m and m = 25.0 kg is laying so that d = 1.00 m of the board hangs off the roof of a building. A con4. A ball m = 0.650 kg and R = 0.170 m is rolling without slipping at a constant speed ve = 3.560 m/s for a time t = 4.110 s.
Answer #1

VW = mig - MG Given Information » large board of length (L) = 6.0m 2) Mass of board (M) = 25 kg 8) board is hanging at the roTool box to not flip up and slide off the mof. torque arting at com & at other end must be equal and opposite. Normal force I4 liven Information. m = 0.650 kg R = 0.1tom Vo = 9.650 m/s t = 4.110s. As shown in above diagram, consider a ball rolling wiGiven Information m = 1230 kg R = 0.230m initial velocity u = 0 m/s angular acceleration (x) = g.11 rad/5² time (t) = 5.345.